integral fraction numerator dx over denominator x square root of x to the power of 4 minus 16 end root end fraction


1
Expert's answer
2021-08-26T18:17:42-0400

"\int\frac{dx}{x\sqrt{x^4-16}}\\
put \space x^4-16=t^2\\
4x^3dx=2tdt\\
\int\frac{dx}{x\sqrt{x^4-16}}\\
=\frac{1}{2}\int\frac{dt}{t^2+16}\\
=\frac{1}{32}\int\frac{dt}{(\frac{t}{4})^2+1}\\
=\frac{4}{32}tan^{-1}(\frac{t}{4})+c\\
=\frac{1}{8}tan^{-1}(\frac{\sqrt{x^4-16}
}{4})+c\\"